OpenStudy (anonymous):

a right cone has a surface area of pi. the radius is 8. white and solve the formula to find the slant height

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

1/8 what?

OpenStudy (anonymous):

is the slant height

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Okay. The area is pi*r*l So the area here is pi So pi = pi*r*l You cancel pi. So 1 = 1*8*l You solve for l and you get 1/8

OpenStudy (anonymous):

l is the slant height